dispensary Borrowed $360,000 on January 2 ,2016, by issuing a 10% serial bond payable that must be paid in three equal annual installments plus interest for the year. The first payment of principal and interest comes due January 2 , 2017. Complete the missing information. Assume bonds are issued at face value. ( for accounts with a $0 balance make sure to enter "0" in the appropriate cell)

2016 2017 2018
Current Liab
Bonds payable
Interest payable
Long term Liab
Bonds payable
Answer #1
Amount $
2016 2017 2018
Current Liab
Bonds payable              120,000              120,000              120,000
Interest payable                 36,000                 24,000                 12,000
Long term Liab
Bonds payable              240,000              120,000                        -  
Workings:
2016 2017 2018
Current Liab
Bonds payable =360000/3 =360000/3 =360000/3
Interest payable =360000*10% =240000*10% =120000*10%
Long term Liab
Bonds payable =360000-120000 =240000-120000 =120000-120000